# MigGPT: Harnessing Large Language Models for Automated Migration of Out-of-Tree Linux Kernel Patches Across Versions

### v1

@@ 5.4 Retrieval Augmentation Module
The retrieval augmentation module is designed to address challenge 1 and challenge 3 encountered during the migration update of out-of-tree kernel patches by LLMs. In challenge 1, LLMs are prone to be misled by similar function signatures when processing function definitions in code snippets, leading to incorrect retrieval of $s_{\text{new}}$ in $\text{file}_{\text{new}}$, which ultimately results in erroneous migrated $s^{\prime}_{\text{new}}$. To overcome this challenge, it is necessary to constrain the LLM’s attention to the target code snippet. As illustrated in Figure 2, the retrieval augmentation module achieves this by constructing a code fingerprint structure ($\text{CFP}_{\text{old}}$) for the old version of the Linux kernel code snippets $s_{\text{old}}$. By analyzing $\text{CFP}_{\text{old}}$, the module extracts the function signatures of the function definitions contained within $s_{\text{old}}$. These function signatures are then used to build a prompt to describe the namespace (“Namespace Prompt”), which is incorporated into the input fed to the LLM. An example is provided in the App. D.2 Figure 8.
On the other hand, challenge 3 highlights that during the migration update of out-of-tree kernel patches by LLMs, there is an issue with missing associated functions. For the LLM’s temporary retrieval result $s_{\text{tmp}}$, we utilize the code fingerprint structures $\text{CFP}_{\text{tmp}}$ and $\text{CFP}_{\text{old}}$ of $s_{\text{tmp}}$ and $s_{\text{old}}$, respectively, and extract from them the sets of internally called associated functions, denoted as $\mathcal{F}_{\text{tmp}}$ and $\mathcal{F}_{\text{old}}$. Then, using string matching techniques, we retrieve from $\text{file}_{\text{new}}$ the code snippets corresponding to the associated function calls $Funccall|$ that satisfy $Funccall|\in\mathcal{F}_{\text{tmp}}\setminus\mathcal{F}_{\text{old}}$. Ultimately, these associated function code snippets are combined with $s_{\text{tmp}}$ to form the complete code snippet $s_{\text{new}}$. An example is provided in the App. D.2 Figure 10.
@@ 5.5 Retrieval Alignment Module
The retrieval alignment module is devised to tackle challenge 2, which was encountered during the migration update of out-of-tree kernel patches by LLMs. Challenge 2 indicates that when the LLM retrieves the target code snippet $s_{\text{new}}$ from the new version of the Linux kernel file $\text{file}_{\text{new}}$, there can be a mismatch between the start and end lines of $s_{\text{new}}$. To address this issue, we need to leverage the information from the first and last lines of the old version code snippet $s_{\text{old}}$ to aid in the localization during the retrieval of $s_{\text{new}}$. As illustrated in Figure 2, we utilize the code fingerprint structure $\text{CFP}_{\text{old}}$ of $s_{\text{old}}$. By taking advantage of its linear structure, we obtain the CFP statements for the first and last lines. These statements are used to construct an “Alignment Prompt”, which describes the information of the first and last lines and is included as part of the input to the LLM. This prompt guides the LLM in performing the retrieval task better by accurately identifying the boundaries of the target code snippet.
Algorithm 1 Retrieval of the target code snippet $s_{\text{new}}$
2:  Output: $s_{\text{new}}$
3:  Generating $\text{CFP}_{\text{old}}$ form $s_{\text{old}}$
4:  Preparing $RetrievalTaskPrompt$
5:  Preparing $ExpertPersonaPrompt$
6:  $\mathcal{S}\leftarrow\operatorname{Extractsignature}(\text{CFP}_{\text{old}})$
7:  $NamespacePrompt\leftarrow\operatorname{Prompt}(\mathcal{S})$
8:  $\mathcal{A}\leftarrow\operatorname{Extractanchor}(\text{CFP}_{\text{old}})$
9:  $AlignmentPrompt\leftarrow\operatorname{Prompt}(\mathcal{A})$
10:  $RetrievalPrompt\leftarrow$
11:   $+RetrievalTaskPrompt+NamespacePrompt$
12:   $+AlignmentPrompt+ExpertPersonaPrompt$
13:  while $q<=m$ do
15:     Generating $\text{CFP}_{\text{tmp}}$ from $s_{tmp}$
16:     if $\operatorname{find}(\mathcal{S},s_{tmp})$ then
17:        break
18:     end if
19:     $q\leftarrow q+1$
20:  end while
21:  $\mathcal{F}_{\text{old}}\leftarrow\operatorname{Funccall}(\text{CFP}_{\text{old}})$
22:  $\mathcal{F}_{\text{tmp}}\leftarrow\operatorname{Funccall}(\text{CFP}_{\text{tmp}})$
23:  $s_{\text{new}}\leftarrow s_{\text{tmp}}+\operatorname{FindCode}(\mathcal{F}_{\text{tmp}}\setminus\mathcal{F}_{\text{new}},\text{file}_{\text{new}})$
24:  return $s_{\text{new}}$
Algorithm 2 Migration of code snippet $s^{\prime}_{\text{new}}$
2:  Output: $s^{\prime}_{\text{new}}$
4:  Preparing $MigrationTaskPrompt$
5:  Preparing $ExpertPersonaPrompt$
6:  $\mathcal{P}\leftarrow\operatorname{PinpointMigrationLocation}(\text{CFP}_{\text{old}},\text{CFP}^{\prime}_{\text{old}})$
7:  $LocationPrompt\leftarrow\operatorname{Prompt}(\mathcal{P})$
8:  $MigrationPrompt\leftarrow+MigrationTaskPrompt$
9:   $+LocationPrompt+ExpertPersonaPrompt$
10:  $s^{\prime}_{\text{new}}\leftarrow\operatorname{LLM}(LocationPrompt,s_{\text{old}},s^{\prime}_{\text{old}},s_{\text{new}})$
11:  return $s^{\prime}_{\text{new}}$
@@ 5.7 Implementation
With the critical code information provided by CFP, we can leverage the Retrieval Augmentation Module and the Retrieval Alignment Module to assist LLMs in more effectively identifying target kernel code snippets $s_{\text{new}}$. Subsequently, with the aid of the Migration Augmentation Module, we facilitate the migration to generate the final code snippet $s^{\prime}_{\text{new}}$. All the prompts of MigGPT are provided in App. D.4.
As illustrated in Algorithm 1, we first need to retrieve the target code snippet $s_{\text{new}}$ from $\text{file}_{new}$. Specifically, using the information contained within $\text{CFP}_{old}$, we can extract a set of critical function signatures $\mathcal{S}$ and a set of key anchor statements $\mathcal{A}$. With this information, we construct the $NamespacePrompt$ and $AlignmentPrompt$, ultimately forming the complete $RetrievalPrompt$. We then query LLMs using the $RetrievalPrompt$ to obtain an initial result $s_{\text{tmp}}$. We check if $s_{\text{tmp}}$ contains items from the target function signature set $\mathcal{S}$. If not, we repeatedly query the LLMs using the $RetrievalPrompt$ (up to $m$ times). If it does contain items from $\mathcal{S}$, we use $\text{CFP}_{old}$ and $\text{CFP}_{tmp}$ to extract newly appeared called functions within $s_{\text{tmp}}$ and retrieve the code snippets where these called functions are defined from $\text{file}_{new}$ as additional supplementary information for $s_{\text{tmp}}$. Finally, we concatenate these two parts of the code snippets to obtain $s_{\text{new}}$.
After obtaining $s_{\text{new}}$, we proceed to migrate it to achieve $s^{\prime}_{\text{new}}$. As shown in Algorithm 2, we utilize the differences between $\text{CFP}_{old}$ and $\text{CFP}^{\prime}_{old}$ to extract the number and positions of migration points and generate the $LocationPrompt$. Further, we formulate the $MigrationPrompt$ and query the LLM to obtain the migrated out-of-tree kernel patch code snippet $s^{\prime}_{\text{new}}$.
@@ 6 Evaluation
In this section, we assess the performance of MigGPT, focusing on the following three research questions:
RQ1 (Performance): How does the performance of MigGPT compare with that of vanilla LLM models?
RQ2 (Ablation): How does each module within MigGPT contribute to the overall migration performance?
@@
LLM Method Type 1 (80) Type 2 (55) All (135) Average Best Match Semantic Match Human Match Best Match Semantic Match Human Match Best Match Semantic Match Human Match Query Times GPT-3.5 Vanilla 20.00% 33.75% 26.25% 20.00% 25.45% 27.27% 20.00% 30.37% 26.67% 1.00 MigGPT 68.75% 68.75% 71.25% 61.82% 54.55% 70.91% 65.93% 62.96% 71.11% 1.28 GPT-4-turbo Vanilla 60.00% 67.50% 65.00% 69.09% 76.36% 78.18% 63.70% 71.11 70.37% 1.00 MigGPT 91.25% 95.00% 96.25% 81.82% 83.64% 89.09% 87.41% 90.37% 93.33% 1.16 DeepSeek-V2.5 Vanilla 61.25% 66.25% 62.50% 65.45% 69.09% 67.27% 62.96% 67.40% 64.44% 1.00 MigGPT 95.00% 97.5% 96.25% 87.27% 90.90% 90.90% 91.85% 94.81% 94.07% 1.22 DeepSeek-V3 Vanilla 68.75% 71.25% 72.50% 74.55% 78.18% 78.18% 71.11% 74.07% 74.81% 1.00 MigGPT 92.50% 93.75% 95.00% 85.45% 78.18% 89.09% 89.63% 87.41% 92.59% 1.22 Llama-3.1-70B Vanilla 58.75% 65.00% 63.75% 61.82% 72.73% 75.55% 60.00% 68.15% 68.15% 1.00 MigGPT 91.25% 92.50% 93.75% 80.00% 81.82% 81.82% 86.67% 88.15% 88.89% 1.24 Average Vanilla 53.75% 60.75% 58.00% 58.18% 64.36% 65.09% 55.56% 62.22% 60.89% 1.00 MigGPT 87.75% 89.50% 90.50% 79.27% 77.82% 84.36% 84.30% 84.74% 88.00% 1.23 $\uparrow$ +34.00% +28.75% +32.50% +21.09% +13.45% +19.27% +28.74% +22.52% +27.11% -
@@
LLM Method Type 1 (80) Type 2 (55) All (135) Best Match Semantic Match Human Match Best Match Semantic Match Human Match Best Match Semantic Match Human Match GPT-3.5 Vanilla 7.50% 5.00% 8.75% 3.64% 3.64% 5.45% 5.93% 4.44% 7.41% MigGPT 37.50% 46.26% 47.50% 38.18% 41.82% 61.82% 37.78% 44.44% 53.33% GPT-4-turbo Vanilla 15.00% 12.50% 18.75% 10.91% 30.91% 23.64% 13.33% 20.00% 20.74% MigGPT 68.75% 82.50% 85.00% 54.55% 76.36 69.09% 62.96% 80.00% 78.52% DeepSeek-V2.5 Vanilla 11.25% 16.25% 18.75% 9.09% 27.27% 21.82% 10.37% 20.74% 20.00% MigGPT 67.50% 80.00% 80.00% 56.36% 74.55% 70.91% 62.96% 77.78% 76.30% DeepSeek-V3 Vanilla 23.75% 37.50% 32.50% 34.55% 54.55% 49.09% 28.15% 44.44% 39.26% MigGPT 81.25% 88.75% 87.50% 65.45% 78.18% 74.55% 74.81% 84.44% 82.22% Llama-3.1-70B Vanilla 3.75% 16.25% 18.75% 7.27% 27.27% 23.64% 5.19% 20.74% 20.74% MigGPT 62.50% 80.00% 81.25% 47.27% 67.27% 72.73% 56.30% 74.81% 77.78% Average Vanilla 12.25% 17.50% 19.50% 13.09% 28.73% 24.73% 12.59% 22.07% 21.63% MigGPT 63.50% 75.50% 76.25% 52.36% 67.64% 69.82% 58.96% 72.30% 73.63% $\uparrow$ +51.25% +58.00% +56.75% +39.27% +38.91% +45.09% +46.37% +50.22% +52.00%
6.1 Evaluation Settings
We assess MigGPT using two benchmarks: the out-of-tree kernel patch migration benchmark from Section 4 and FixMorph’s CVE patch backporting benchmark (Shariffdeen et al., 2021a), which includes 350 instances. For baselines, we use vanilla LLMs, including GPT-3.5 (OpenAI, 2023a), GPT-4-turbo (OpenAI, 2023b), DeepSeek-V2.5 (DeepSeek-AI et al., 2024a), DeepSeek-V3 (DeepSeek-AI et al., 2024b), and Llama-3.1-70B-Instruct (Dubey et al., 2024), as they are widely recognized for their advanced capabilities and performance in code-related tasks, along with patch backporting tools like Patch (GNU, 2020), SyDIT (Meng et al., 2011), and FixMorph (Shariffdeen et al., 2021a). Evaluation metrics include “best match” (exact code similarity after removing spaces, line breaks, and tab characters), “semantic match” (CodeBLEU with a 0.9 threshold for binary classification, detailed in App. E.2(Ren et al., 2020), and “human match” (developer-judged functional equivalence). The hyperparameter $m$ is set to 3.
For each migration sample $(s_{\text{old}},s^{\prime}_{\text{old}},\text{file}_{\text{new}},\hat{s}_{\text{new}},\hat{s}^{\prime}_{\text{new}})$ in our benchmark, we evaluate vanilla LLMs using two distinct strategies: One-step Strategy: The LLM directly generates the migrated code snippet $s^{\prime}_{\text{new}}$ by taking the triplet $(s_{\text{old}},s^{\prime}_{\text{old}},\text{file}_{\text{new}})$ as input. Two-step Strategy: The process is divided into two phases. First, the LLM identifies the corresponding new version code snippet $s_{\text{new}}$ using the pair $(s_{\text{old}},\text{file}_{\text{new}})$. Then, the LLM generates $s^{\prime}_{\text{new}}$ by taking the triplet $(s_{\text{old}},s^{\prime}_{\text{old}},s_{\text{new}})$ as input.
Figure 5: The semantic match accuracy of target code snippets retrieval task and target code snippets migration task across various LLMs. “One-step” indicates the direct utilization of an LLM to complete the migration task in a single step.

### v2

@@ 5.4 Retrieval Augmentation Module
The retrieval augmentation module is designed to address challenge 1 and challenge 3 encountered during the migration update of out-of-tree kernel patches by LLMs. In challenge 1, LLMs are prone to be misled by similar function structure when processing function definitions in code snippets, leading to incorrect retrieval of $s_{\text{new}}$ in $\text{file}_{\text{new}}$, which ultimately results in erroneous migrated $s^{\prime}_{\text{new}}$. To overcome this challenge, it is necessary to constrain the LLM’s attention to the target code snippet. As illustrated in Figure 2, the retrieval augmentation module achieves this by constructing a code fingerprint structure ($\text{CFP}_{\text{old}}$) for the old version of the Linux kernel code snippets $s_{\text{old}}$. By analyzing $\text{CFP}_{\text{old}}$, the module extracts the function signatures of the function definitions contained within $s_{\text{old}}$. These function signatures are then used to build a prompt to describe the structure information (“Structure Prompt"), which is incorporated into the input fed to the LLM. An example is provided in the App. D.2 Figure 8.
[…]
@@ 5.5 Retrieval Alignment Module
The retrieval alignment module is devised to tackle challenge 2, which was encountered during the migration update of out-of-tree kernel patches by LLMs. Challenge 2 indicates that when the LLM retrieves the target code snippet $s_{\text{new}}$ from the new version of the Linux kernel file $\text{file}_{\text{new}}$, there can be a mismatch between the boundary line of $s_{\text{new}}$. To address this issue, we need to leverage the information from the first and last lines of the old version code snippet $s_{\text{old}}$ to aid in the localization during the retrieval of $s_{\text{new}}$. As illustrated in Figure 2, we utilize the code fingerprint structure $\text{CFP}_{\text{old}}$ of $s_{\text{old}}$. By taking advantage of its linear structure, we obtain the CFP statements for the first and last lines. These statements are used to construct an “Alignment Prompt", which describes the information of the first and last lines and is included as part of the input to the LLM. This prompt guides the LLM in performing the retrieval task better by accurately identifying the boundaries of the code snippet.
@@ 5.7 Implementation
With the critical code information provided by CFP, we can leverage the Retrieval Augmentation Module and the Retrieval Alignment Module to assist LLMs in more effectively identifying target kernel code snippets $s_{\text{new}}$. Subsequently, with the aid of the Migration Augmentation Module, we facilitate the migration to generate the final code snippet $s^{\prime}_{\text{new}}$. All the prompts and the algorithm of MigGPT are provided in App. D.4.
As illustrated in Figure 2, we first need to retrieve code snippet $s_{\text{new}}$ from $\text{file}_{new}$. Specifically, using the information contained within $\text{CFP}_{old}$, we can extract a set of critical function signatures $\mathcal{S}$ and a set of key anchor statements $\mathcal{A}$. With this information, we construct the $StructurePrompt$ and $AlignmentPrompt$, ultimately forming the complete $RetrievalPrompt$. We then query LLMs using the $RetrievalPrompt$ to obtain an initial result $s_{\text{tmp}}$. We check if $s_{\text{tmp}}$ contains items from the target function signature set $\mathcal{S}$. If not, we repeatedly query the LLMs using the $RetrievalPrompt$ (up to $m$ times). If it does contain items from $\mathcal{S}$, we use $\text{CFP}_{old}$ and $\text{CFP}_{tmp}$ to extract newly appeared called functions within $s_{\text{tmp}}$ and retrieve the code snippets where these called functions are defined from $\text{file}_{new}$ as additional supplementary information for $s_{\text{tmp}}$. Finally, we concatenate these two parts of the code snippets to obtain $s_{\text{new}}$. After obtaining $s_{\text{new}}$, we proceed to migrate it to achieve $s^{\prime}_{\text{new}}$. We utilize the differences between $\text{CFP}_{old}$ and $\text{CFP}^{\prime}_{old}$ to extract the number and positions of migration points and generate the $LocationPrompt$. Further, we formulate the $MigrationPrompt$ and query the LLM to obtain the migrated out-of-tree kernel patch code snippet $s^{\prime}_{\text{new}}$.
@@ 6 Evaluation
In this section, we assess the performance of MigGPT, focusing on the following questions:
RQ1 (Performance): How does the performance of MigGPT compare with that of vanilla LLM?
RQ2 (Ablation): How does each module within MigGPT contribute to the overall performance?
@@
6.1 Evaluation Settings
We assess MigGPT using two benchmarks: the out-of-tree kernel patch migration benchmark from Section 4 and FixMorph’s CVE patch backporting benchmark FixMorph , which includes 350 instances. For baselines, we use vanilla LLMs, including GPT-3.5 gpt-3.5 , GPT-4-turbo gpt4 , OpenAI-o1 openai-o1 , DeepSeek-V2.5 deepseek-2.5 , DeepSeek-V3 deepseek-3 , Deepseek-R1 deepseek-r1 , Llama-3.1-8B and Llama-3.1-70B-Instruct llama-3.1 , as they are widely recognized for their advanced performance, along with previous migration efforts like FixMorph FixMorph , TSBPORT TSBPORT and PPatHF PPatHF . Evaluation metrics include “best match" (exact code similarity after removing spaces, line breaks, and tab characters), “semantic match" (CodeBLEU with a 0.9 threshold for binary classification, detailed in App. F.5codebleu , and “human match" (developer-judged functional equivalence, detailed in App. F.2). We also considered the compilation success rate, as detailed in App. F.3. The hyperparameter $m$ is set to 3.
For each sample $(s_{\text{old}},s^{\prime}_{\text{old}},\text{file}_{\text{new}},\hat{s}_{\text{new}},\hat{s}^{\prime}_{\text{new}})$ in our benchmark, we evaluate vanilla LLMs using two distinct strategies: One-step Strategy: The LLM directly generates the migrated code snippet $s^{\prime}_{\text{new}}$ by taking the triplet $(s_{\text{old}},s^{\prime}_{\text{old}},\text{file}_{\text{new}})$ as input. Two-step Strategy: The process is divided into two phases. First, the LLM identifies the corresponding new version code snippet $s_{\text{new}}$ using the pair $(s_{\text{old}},\text{file}_{\text{new}})$. Then, the LLM generates $s^{\prime}_{\text{new}}$ by taking the triplet $(s_{\text{old}},s^{\prime}_{\text{old}},s_{\text{new}})$ as input.
@@ 6.1 Evaluation Settings
LLM Method Type 1 (80) Type 2 (55) All (135) Average Best Match Semantic Match Human Match Best Match Semantic Match Human Match Best Match Semantic Match Human Match Query Times GPT-3.5 Vanilla 20.00% 33.75% 26.25% 20.00% 25.45% 27.27% 20.00% 30.37% 26.67% 1.00 MigGPT 68.75% 68.75% 71.25% 61.82% 54.55% 70.91% 65.93% 62.96% 71.11% 1.28 GPT-4-turbo Vanilla 60.00% 67.50% 65.00% 69.09% 76.36% 78.18% 63.70% 71.11% 70.37% 1.00 MigGPT 91.25% 95.00% 96.25% 81.82% 83.64% 89.09% 87.41% 90.37% 93.33% 1.16 OpenAI-o1 Vanilla 76.25% 82.50% 80.00% 81.82% 85.45% 92.73% 78.52% 83.70% 85.19% 1.00 MigGPT 96.25% 97.50% 96.25% 85.45% 89.09% 92.73% 91.85% 94.07% 94.81% 1.25 DeepSeek-V3 Vanilla 68.75% 71.25% 72.50% 74.55% 78.18% 78.18% 71.11% 74.07% 74.81% 1.00 MigGPT 92.50% 93.75% 95.00% 85.45% 78.18% 89.09% 89.63% 87.41% 92.59% 1.22 DeepSeek-R1 Vanilla 72.50% 76.25% 77.50% 63.64% 70.91% 74.55% 68.89% 74.07% 76.30% 1.00 MigGPT 95.00% 95.00% 95.00% 80.00% 85.45% 87.27% 88.89% 91.11% 91.85% 1.23 Llama-3.1-8B Vanilla 37.50% 46.25% 43.75% 43.64% 47.27% 45.45% 40.00% 46.67% 44.44% 1.00 MigGPT 77.50% 80.00% 81.25% 70.91% 74.55% 78.18% 74.81% 77.78% 80.00% 1.36 Llama-3.1-70B Vanilla 58.75% 65.00% 63.75% 61.82% 72.73% 75.55% 60.00% 68.15% 68.15% 1.00 MigGPT 91.25% 92.50% 93.75% 80.00% 81.82% 81.82% 86.67% 88.15% 88.89% 1.29 Average Vanilla 56.25% 63.26% 61.25% 59.22% 65.19% 67.42% 57.46% 64.02% 63.70% 1.00 MigGPT 87.50% 88.93% 76.25% 77.92% 78.18% 84.16% 83.60% 84.55% 87.51% 1.26 $\uparrow$ +31.25% +25.67% +15.00% +18.70% +12.99% +16.74% +26.14% +20.53% +23.81% -
@@
LLM Method Type 1 (80) Type 2 (55) All (135) Best Match Semantic Match Human Match Best Match Semantic Match Human Match Best Match Semantic Match Human Match GPT-3.5 Vanilla 7.50% 5.00% 8.75% 3.64% 3.64% 5.45% 5.93% 4.44% 7.41% MigGPT 37.50% 46.26% 47.50% 38.18% 41.82% 61.82% 37.78% 44.44% 53.33% GPT-4-turbo Vanilla 15.00% 12.50% 18.75% 10.91% 30.91% 23.64% 13.33% 20.00% 20.74% MigGPT 68.75% 82.50% 85.00% 54.55% 76.36% 69.09% 62.96% 80.00% 78.52% OpenAI-o1 Vanilla 20.00% 28.75% 30.00% 18.18% 30.91% 27.27% 19.26% 29.63% 28.89% MigGPT 77.50% 90.00% 90.00% 60.00% 76.36% 74.55% 70.37% 84.44% 83.70% DeepSeek-V3 Vanilla 23.75% 37.50% 32.50% 34.55% 54.55% 49.09% 28.15% 44.44% 39.26% MigGPT 81.25% 88.75% 87.50% 65.45% 78.18% 74.55% 74.81% 84.44% 82.22% DeepSeek-R1 Vanilla 53.75% 60.00% 62.50% 40.00% 54.55% 56.36% 48.15% 57.78% 60.00% MigGPT 72.50% 85.00% 81.25% 69.09% 81.82% 78.18% 71.11% 83.70% 80.00% Llama-3.1-8B Vanilla 5.00% 12.50% 16.25% 0% 20.00% 25.45% 2.96% 15.56% 20.00% MigGPT 36.25% 65.00% 67.50% 25.45% 52.73% 56.36% 31.85% 60.00% 62.96% Llama-3.1-70B Vanilla 3.75% 16.25% 18.75% 7.27% 27.27% 23.64% 5.19% 20.74% 20.74% MigGPT 62.50% 80.00% 81.25% 47.27% 67.27% 72.73% 56.30% 74.81% 77.78% Average Vanilla 18.39% 24.64% 26.79% 16.36% 31.69% 30.13% 17.57% 27.51% 28.15% MigGPT 62.32% 76.78% 77.14% 51.43% 67.79% 69.61% 57.88% 73.12% 74.07% $\uparrow$ +43.93% +52.14% +50.36% +35.06% +36.10% +39.48% +40.32% +45.61% +45.92%
Figure 5: The semantic match accuracy of target code snippets retrieval task and target code snippets migration task across various LLMs. “One-step" indicates the direct utilization of an LLM to complete the migration task in a single step.